Please help me to solve this question
A watch gains 30 minutes every day. If it is set as 6:00 AM. How many minutes it gains at 10:00 PM?

Answer by checkley79(3341)

30 MINUTES = .5 HOURS.
30/24=X/4 CROSS MULTIPLY.
24X=4*30
24X=120
X=120/24
X=5 MINUTES ANS.
PROOF:
30/24=5/4
24*5=30*4
120=120